0 0
Add a comment Improve this question Transcribed image text
Answer #1

Itl 2 1t 지 ) もーう Room Raure it is seen ttrad both *)&XGa) are erev en nahure e aah stheC) incr and the banducth ralueao Sin xclear all t--10: . 01: 10; x-exp (-abs (t)) subplot (1,2,1); plot (t, x): title (x(t) ) k-0: for w--100:1:100 X(k)trapz (t,clear all t -10:.01:10: x-exp (-abs (t)) subplot (1,2,1); plot (t,x); title(x(t) ) k=0; for w=-100: 1 : 100 k=k+1; x (k) =tt -10:.01:10: T-input (enter T n-t./T x-exp (-abs (t) subplot (1,3,1); plot (t,x); title(x(t) ) h-sin (a) ./a; subplot (1x(t) 20 0.9 0.8 15 0.7 0.6 10 0.5 0.4 바 0.3 0.2 0.1 10 6 4 0 2 46 8 100 -80 -60 40 20 020 40 60 80 100xL(t) for wm 5 xL(t) for wm-2 1.6 1.4 1.2 0.9 0.8 0.7 0.6 0.5 0.4 0.3 0.2 0.1 0.7 0.6 0.5 0.4 0.3 0.2 0.8 0.6 0.4 0.2 0.2 -0.x(t) h(t) x1(t) for pi/2 0.03 0.9 0.8 0.7 0.6 0.5 0.4 0.3 0.2 0.1 0.8 0.025 0.6 0.02 0.4 0.015 0.2 0.01 0.005 0.2 0 10-0 00 1x(t) h(t) x1(t) for pi/10 0.9 0.8 0.7 0.6 0.5 0.4 0.3 0.2 0.1 0.01 -0.02 0.03 0.04 0.05 0.06 0.07 -0.08 -0.09 10 20 0.8 0.6 0